An economist makes an assumption that each additional year of education causes future wages to rise by 7 percent. In this​ model, if a person with 12 years of education makes ​$21 000 per​ year, then a person with 4​-year college degree would earn ​? per year.​ (Round your intermediate calculations to two decimal​ places.)
Answer #1

Answer : Additional education increase the wages by 7% per year educational year.

After 12 years education the wage is $21,000 per year.

College degree requires 4 year and each year makes the 7% increase in wages. So,

For 1st year of college degree the wage rate increase = 7% of $21,000 + $21,000 = (0.07 * 21,000) + 21,000 = $1,470 + 21,000 = $22,470.

For 2nd year of college degree the wage rate increase = 7% of $22,470 + $22,470 = (0.07 * 22,470) + 22470 = 1,572.90 + 22,470 = $24,042.90 .

For 3rd year of college degree the wage rate increase = 7% of $24,042.90 + $24,042.90 = (0.07 * 24,042.90) + 24,042.90 = 1,683.00 + 24,042.90 = $25,725.90

For 4th year of college degree the wage rate increase = 7% of $25,725.90 + $25,725.90 = 1,800.81 + 25,725.90 = $27,526.71

Therefore, after college degree a person will earn $27,526.71 per year.